Hi
The distance between town A and town B is 35km. A cyclist leaves town A for town B at a speed of 24km per hour at 11.55am. At the same time a pedestrian leaves town B for town A at 6km per hour. At what distance from A will they meet each other.
Thanks

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

The distance between town A and town B is 35km.
A cyclist leaves town A for town B at a speed of 24km per hour at 11.55am.
At the same time a pedestrian leaves town B for town A at 6km per hour.
At what distance from A will they meet each other.
--------------
The start time is not relevant.
Moving in opposite directions ---> the speeds are added = 30 km/hr
Time = 35/30 = 7/6 hours
Distance at 24 km/hr = 24*(7/6) = 28 km from town A

Answer by greenestamps(13200)



The ratio of the two speeds is 24:6 = 4:1.

The distances are the same, and they start at the same time, so the ratio of the distances they cover before they meet is 4:1.

So they meet when the cyclist has traveled 4/5 of 35km, which is 28km.

ANSWER: They will meet 28km from town A.